# The Statistical Mechanics of Hawking Radiation

### v3

@@ 2.1 Metric of Choice
Particle worldlines near a black hole follow the Schwarzschild metric, written here in Eddington-Finkelstein (EF) coordinates [8, 9] using the metric signature (+,,,)(+,-,-,-)( + , - , - , - ):
@@
where $d\Omega^{2}=d\theta^{2}+\sin^{2}\theta d\phi^{2}$ represents the solid angle part of the metric. In EF coordinates, the modified time coordinate $dt^{\prime}$ is defined as $dt\pm(dr^{*}-dr)$, where the sign depends on the particle’s radial direction: either “ingoing” ($+$) or “outgoing” (--). Additionally, $r^{*}$ is the tortoise coordinate that satisfies $dr^{*}/dr=1/f(r)$ with $f(r)=(1-r_{S}/r)$.
For an evaporating black hole, ingoing EF coordinates are used with the Vaidya metric instead of the Schwarzschild metric [10]. The Vaidya metric, given originally in Ref. [10] in the (,+,+,+)(-,+,+,+)( - , + , + , + ) signature, is revised here for consistency with the (+,,,)(+,-,-,-)( + , - , - , - ) signature:
@@
with $m_{P}\equiv\sqrt{\hbar/G}$ as the Planck mass, $M_{0}$ as the initial black hole mass (which is gauged to be astronomical), and $M$ as the instantaneous mass. The functions $g(z)$ and $h(z)$ are polynomials in $z$, given in Ref. [10] as Eq. (85). If we omit higher-order terms of $z$, given the relevant range $z\in(0,1]$, we define
@@
where $\beta=({2^{14}3^{4}5\pi})^{-1}$, and the parameters $\xi$, $k_{3,4,5,6}$ and $f_{0}$ depend on particle spin. E.g. for photons (spin-1): $\xi=-1248$, $k_{3}=81.80$, $k_{4}=-770.42$, $k_{5}=65.38$, $k_{6}=-942.18$, and $f_{0}=2.4346$ [10, 11, 12]. Other known values of $\xi$ include $168$ for spin-1/2 particles and $20352$ for gravitons (spin-2) [10]; the $k$ parameters are unknown for these cases.
@@
The common leading factor, $6\xi\beta$, evaluates numerically as $4.84\times 10^{-4}$ for spin-1/2 particles, $-3.59\times 10^{-4}$ for photons, and $5.86\times 10^{-3}$ for gravitons. Given the small magnitude of these factors and their proximity to zero, we approximate – however naively – $g(z)$ and $h(z)$ to be zero. This simplifies the metric functions to
$\psi(z)\approx-4\alpha\frac{m_{P}^{2}}{M_{0}^{2}}\ln(z)\quad\mathrm{and}\quad\mu(z)\approx M,$ (7)
@@ 2.2 Worldline Lagrangian
This form reveals that the Lagrangian consists of the relativistic energy of the emitted particle and a quantum term, suggesting a discrete energy structure reminiscent of Bohr’s hydrogen model. If we model black holes as structured potential wells containing $s$ total quantum states, the availability and abundance of these states depend on the black hole’s size. Since the event horizon encompasses $N$ quantum areas with $A_{Q}=16\pi l_{P}^{2}$ [13, 14], we have the relation:
$N=\frac{A_{\text{BH}}}{A_{{Q}}}\implies\frac{M_{0}^{2}}{m_{P}^{2}},$ (14)
where $M_{0}$ is the initial black hole mass and $m_{P}$ is the Planck mass. Furthermore, the emitted Hawking particle’s energy state $q$ can be interpreted as an available emission state from the black hole surface. For a stationary, spherical black hole with azimuthal symmetry ($\phi=2\pi$), the energy state $q$ becomes the relevant parameter for describing the evaporation rate of black hole surface states:
$\mathcal{L}=m_{H}-2\pi\hbar\frac{dq}{d\tau}.$ (15)
@@ 2.3 A Generalized Liouville Theorem
[…]
@@
Since the black hole horizon is macroscopically a curved, dynamical 4D manifold, the set of accessible microstates evolves in a way that must be consistent with the ergodic hypothesis across all possible foilations [16]. This suggests that the rate of change of $\Omega$ (here in a 8D phase space) should be expressed in terms of a general relativistic continuity equation, incorporating the covariant divergence of microstate flux:
$\frac{d\Omega}{d\Sigma_{0}}\Big{|}_{t-t^{\prime}\simeq d\tau}=\sum_{i=1}^{N}\iint\frac{d^{4}\mathbf{x}d^{4}\mathbf{p}}{(2\pi\hbar)^{4}\sqrt{-g}}\nabla_{\mu}(\rho_{i}u^{\mu})=0.$ (17)
@@
As the black hole evaporates, Hawking radiation leads to a net decrease in the number of surface microstates. To account for this, we introduce a “system-environment ensemble,” where microstates lost from the black hole surface are transferred to the radiation background:
@@
$\frac{d}{d\tau}\left(\iint\rho_{\text{BH}}\frac{d^{4}\mathbf{x}d^{4}\mathbf{p}}{(2\pi\hbar)^{4}}+\iint\rho_{\text{rad}}\frac{d^{4}\mathbf{x}d^{4}\mathbf{p}}{(2\pi\hbar)^{4}}\right)=0.$ (20)
@@ 2.4 State Evaporation
To simplify the 8-dimensional theorem, we decompose the physical and momentum-space volumes into an 6-dimensional subspace and a seperate 2D element associated with the emitted state:
$\frac{dq}{d\tau}\Rightarrow\frac{d}{d\tau}\iint\frac{d^{3}\mathbf{x}d^{3}\mathbf{p}}{V^{3}_{x}V^{3}_{p}}\iint\rho_{\text{BH}}(\sigma,\tau)\frac{d\sigma dp}{2\pi\hbar}.$ (21)
@@
To enforce one-particle occupancy per microstate, the probability distribution must be well-defined within the intervals $\sigma\in(\infty,l]$ (where the radiation field extends infinitely but is cut off at $l$, the instantaneous black hole radius) and $p\in[0,m_{H}]$. We define the microstate density heuristically as a square-area density:
@@
where $\tau_{\mathrm{evap}}$ is the total black hole evaporation time [17, 18], given by
$\tau_{\mathrm{evap}}=\frac{5120\pi G^{2}M^{3}}{\hbar}.$ (23)
@@
$\frac{dq}{d\tau}=\frac{-m_{H}r_{S}}{4\pi\hbar}\int_{\infty}^{l}\frac{d\sigma}{\sigma^{2}}=\frac{1}{4\pi\hbar}\frac{m_{H}\,r_{S}}{l}.$ (24)
Here, the integral’s lower limit at $\sigma\rightarrow\infty$ naturally vanishes due to $1/\infty=0$, ensuring that only the boundary at $l$ contributes to the evaporation rate.
Since the evaporation rate is positive and non-zero, the transfer of states from the black hole to the radiation background is irreversible. This follows from the fact that, under semi-classical assumptions, a quantum state escaping the horizon cannot be reabsorbed by the black hole. Upon instantaneous measurement, the black hole’s instantaneous radius $l$ is equivalent to the Schwarzschild radius $r_{S}$. Substituting this into Eq. (24) and incorporating it into Eq. (15), we define the Lagrangian as a single energy term:
$E=\frac{1}{2}m_{H}.$ (25)
To determine the mass of an emitted Hawking particle, recall our quasi-quantum-classical-correspondence: $m_{H}r^{2}\dot{\phi}=q\hbar$. We can use our conditions of azimuthal summetry ($\phi=2\pi$) and proper time variation in $r$ to yield
@@
implying Eq. (10). Our obtained mass depends on the energy state $q$, whose evaporation rate relies strictly on the one-particle occupancy limit. This reinforces this mass relation as the mass of one Hawking particle, and we may gauge $q=1$ as a result. Conventionally, the Hawking particle mass is obtained by the reduced Compton mass formula, $m=\hbar/\lambda_{C}$. Since Hawking radiation originates at the horizon, the Compton wavelength corresponds to the black hole circumference, $\lambda_{C}=2\pi r_{S}$. This requires $q=1$ in Eq. (26); refer to footnote [^1] We eventually verify that $q=1$ is fundamental by evaluating Eq. (24) using the $q$-independent Compton mass definition of $m_{H}$.. From both directions, we yield
@@
Therefore, the energy of the Hawking particle is exactly the thermal energy of the black hole:
$E_{\mathrm{tot}}=\frac{\hbar}{8\pi GM}\equiv k_{B}T_{BH}.$ (28)
@@ 2.5 Total Energy States $s$
${dq}=\frac{1}{8\pi^{2}}\frac{d\tau}{l}.$ (29)
To relate this to the black hole’s evaporation, we impose the condition $d\tau=dl$, linking the change in proper time to the decrease in the black hole’s instantaneous radius. Since the Hawking particle follows a worldine with $ds=d\tau$, this establishes a direct correspondence between the emitted particle’s trajectory and the shrinking black hole radius. Integrating over the range $l\in[2l_{P},r_{S,0}]$, we obtain
@@
[…]
For astrophysical black holes with a minimum mass of $3M_{\odot}$, a single quantum state is emitted each time a Hawking particle escapes the event horizon. The remaining quantum states on the black hole surface, therefore, correspond to the remaining number of surface particles along the event horizon. This defines the total number of quantum states as $\lfloor s\rfloor=N\lfloor q\rfloor$, where $N$ is given by Eq. (14). Consequently, the total number of states is
$s=\frac{M_{0}^{2}}{8\pi^{2}m_{P}^{2}}\ln\left(\frac{M_{0}}{m_{P}}\right).$ (31)

### v4

@@ 2.1 Metric of Choice
Particle worldlines near a black hole follow the Schwarzschild metric, written here in Eddington-Finkelstein (EF) coordinates [8, 9] using the metric signature (+,,,)(+,-,-,-):
@@
where $d\Omega^{2}=d\theta^{2}+\sin^{2}\theta d\phi^{2}$ represents the solid angle part of the metric. In EF coordinates, the modified time coordinate $dt^{\prime}$ is defined as $dt\pm(dr^{*}-dr)$, where the sign depends on the particle’s radial direction: either “ingoing” (++) or “outgoing” (-). Additionally, $r^{*}$ is the tortoise coordinate that satisfies $dr^{*}/dr=1/f(r)$ with $f(r)=(1-r_{S}/r)$.
For an evaporating black hole, ingoing EF coordinates are used with the Vaidya metric instead of the Schwarzschild metric [12]. The Vaidya metric, given originally in Ref. [12] in the (,+,+,+)(-,+,+,+) signature, is revised here for consistency with the (+,,,)(+,-,-,-) signature:
@@
with $m_{P}\equiv\sqrt{\hbar/G}$ as the Planck mass, $M_{0}$ as the initial black hole mass (which is gauged to be astronomical), and $M$ as the instantaneous mass. We can define the ratio between the squares of the initial black hole and Planck masses as an “initial abundance number”:
$N_{0}=\frac{M_{0}^{2}}{m_{P}^{2}},$ (5)
such that the initial black hole’s surface area $A_{0}=16\pi G^{2}M_{0}^{2}$ is divided into $N_{0}$ quantum black hole surface areas via $A_{Q}=16\pi l_{P}^{2}$ [10, 11], where $l_{P}\equiv\sqrt{G\hbar}$ is the Planck length.
The functions $g(z)$ and $h(z)$ are polynomials in $z$, given in Ref. [12] as Eq. (85). If we omit higher-order terms of $z$, given the relevant range $z\in(0,1]$, we define
@@
where $\beta=({2^{14}3^{4}5\pi})^{-1}$, and the parameters $\xi$, $k_{3,4,5,6}$ and $f_{0}$ depend on particle spin. E.g. for photons (spin-1): $\xi=-1248$, $k_{3}=81.80$, $k_{4}=-770.42$, $k_{5}=65.38$, $k_{6}=-942.18$, and $f_{0}=2.4346$ [12, 13, 14]. Other known values of $\xi$ include $168$ for spinors (spin-1/2 particles) and $20352$ for gravitons (spin-2) [12]; the $f_{0}$ and $k$ parameters are unknown for these cases.
@@
The common leading factor, $6\xi\beta$, evaluates numerically as $4.84\times 10^{-4}$ for spinors, $3.59\times 10^{-4}$ for photons, and $5.86\times 10^{-3}$ for gravitons. Given the small magnitude of these factors and their proximity to zero, we approximate – however naively – $g(z)$ and $h(z)$ to be zero. This simplifies the metric functions to the following, using $N_{0}$ via Eq. (5):
$\psi(z)\approx-\frac{4\alpha}{N_{0}}\ln(z)\quad\mathrm{and}\quad\mu(z)\approx M,$ (8)
@@
Despite proceeding with the Schwarzschild metric, we want to present metric modifications that are anticipated when taking $g(z)$ and $h(z)$ via Eq. (7) into proper account. Expanding out the polynomials and neglecting higher orders of $z$ due to smallness, the metric functions read as
$\begin{split}&\psi(z)\approx-\frac{2}{N_{0}}\left[9\xi\beta+2\alpha\ln(z)\right],\\ &\mu(z)\approx M\left(1+\frac{6\xi\beta}{N_{0}}\right),\end{split}$ (10)
which modifies the Vaidya metric as follows, in terms of $z$:
$\begin{split}&ds^{2}=\mathcal{E}(z)\left[1-z\left(1+\frac{6\xi\beta}{N_{0}}\right)\right]dv^{2}\\ &\quad\quad\quad\quad\quad\quad\quad\quad-2\mathcal{E}(z)^{2}dvdr-r^{2}d\Omega^{2},\\ &\mathcal{E}(z):=\left(e^{9\xi\beta}z^{2\alpha}\right)^{N_{0}/4}.\end{split}$ (11)
The modification’s inclusion of $\xi$ suggests a more explicit treatment of spin-dependent particle types, which is encouraged to pursue for future work. While the reduced forms of the metric functions are mutual for spinors, photons, and gravitons, the analysis on photons should involve the explicit form of the polynomials via Eq. (6), as all parameters are known.
@@ 2.2 Worldline Lagrangian
This form reveals that the Lagrangian consists of the relativistic energy of the emitted particle and a quantum term, suggesting a discrete energy structure reminiscent of Bohr’s hydrogen model. If we model black holes as structured potential wells containing $s$ total quantum states, the availability and abundance of these states depend on the black hole’s size by decomposing the black hole surface into $N$ quantum areas, as initially defined as Eq. (5) for an initial mass $M_{0}$ Thus, for an instantaneous mass, $N=M^{2}/m_{P}^{2}$.
The emitted Hawking particle’s energy state $q$ can be interpreted as an available emission state from the black hole surface. Therefore, the state $q$ has an associating emission rate $dq/d\tau$ to describe evaporation from the spherical horizon. This is paired with the rate $d\phi/d\tau$ via a total derivative on the coupled term $q\phi$, factoring in the reduced Planck constant:
$\frac{d(q\hbar\,\phi)}{d\tau}=q\hbar\frac{d\phi}{d\tau}+\phi\hbar\frac{dq}{d\tau}.$ (17)
Utilizing $q\hbar=L$ in the total derivative term (the left-hand side of Eq. [17]) and enforcing Eq. (15) defines a torque-like quantity $d(L\phi)/d\tau$. We can therefore define $q\hbar\,d\phi/d\tau$ in Eq. (16) by re-arranging Eq. (17). This redefinition is specifically useful when considering a stationary, spherical black hole with azimuthal symmetry (i.e., $d\phi=0$ in Eq. [16]). This reveals that $\phi=\text{const.}$ and the Lagrangian would only have the relativistic energy term with no quantum contribution. Redefining $d\phi/d\tau$ via Eq. (17) not only preserves the quantum contribution via the state emission rate $dq/d\tau$, but it also introduces a torque contribution $dL/d\tau$ via Eq. (15) under a constant $\phi$. Constraining $\phi=\text{const.}$ involves, effectively, integrating over all azimuthal angles to gauge $\phi=2\pi$. This yields:
$\mathcal{L}=m_{H}+2\pi\hbar\frac{dq}{d\tau}-2\pi\frac{dL}{d\tau}.$ (18)
@@ 2.3 A Generalized Liouville Theorem
In Eq. (18), the rate $dq/d\tau$ describes the evaporation of a quantum state from the black hole, transitioning into a corresponding quantum state in the radiation background. This process parallels Liouville’s theorem, which states that in a closed Hamiltonian system, the total number of microstates $\Omega$ remains constant if the density of a 2nn-dimesnional phase space is constant [15]:
@@
Eq. (19) is applicible to classical Hamiltonian systems, typically for $n=3$. To generalize this for relativistic Hamiltonian systems, the set of accessible microstates evolves in a way that must be consistent with the ergodic hypothesis across all possible foilations [16]. This suggests that the rate of change of $\Omega$ (here in a 8D phase space, i.e. $n=4$) should be expressed in terms of a general relativistic continuity equation, incorporating the covariant divergence of microstate flux:
$\frac{d\Omega}{d\Sigma_{0}}\Big{|}_{t-t^{\prime}\simeq d\tau}=\sum_{i=1}^{N_{0}}\iint\frac{d^{4}\mathbf{x}d^{4}\mathbf{p}}{(2\pi\hbar)^{4}\sqrt{-g}}\nabla_{\mu}(\rho_{i}u^{\mu})=0.$ (20)
@@
To apply this specifically to black hole evaporatation, Hawking radiation leads to a net decrease in the number of surface microstates. This is given by the the conversion of the initial number $N_{0}$ via Eq. (5) into the instantaneous number $N$, such that $N\in[1,N_{0}]$ and $N_{0}-N$ quantifies the number of quantum areas already radiated from the black hole surface. To account for this, we introduce a “system-environment ensemble,” where microstates lost from the black hole surface are transferred to the radiation background:
@@
$\frac{d}{d\tau}\iint\rho_{\text{BH}}\frac{d^{4}\mathbf{x}d^{4}\mathbf{p}}{(2\pi\hbar)^{4}}+\frac{d}{d\tau}\iint\rho_{\text{rad}}\frac{d^{4}\mathbf{x}d^{4}\mathbf{p}}{(2\pi\hbar)^{4}}=0.$ (23)
@@ 2.4 State Evaporation
Considering the black hole state density contribution in Eq. (23) as the state evaporation rate $dq/d\tau$, we simplify the 8-dimensional theorem by decomposing the physical and momentum-space volumes into an 6-dimensional subspace and a seperate 2D element associated with the emitted state:
$\frac{dq}{d\tau}=\frac{d}{d\tau}\iint\frac{d^{3}\mathbf{x}d^{3}\mathbf{p}}{V^{3}_{x}V^{3}_{p}}\iint\rho_{\text{BH}}(\sigma,\tau)\frac{d\sigma dp}{2\pi\hbar}.$ (24)
@@
To enforce one-particle occupancy per microstate, the probability distribution must be well-defined within the intervals $\sigma\in[l,\infty)$ (where the radiation field radially extends out to infinity from a minimum cut off at $l$, the instantaneous black hole radius) and $p\in[0,m_{H}]$. We define the microstate density heuristically as a square-area density:
@@
such that it exhibits an inverse square-law for radiation intensity. Also, $\tau_{\mathrm{evap}}$ is the total black hole evaporation time [17, 18], given by
$\tau_{\mathrm{evap}}=\frac{5120\pi G^{2}M_{0}^{3}}{\hbar}$ (26)
for the initial black hole mass.
@@
$\begin{split}\frac{dq}{d\tau}&=\frac{d}{d\tau}\left((\tau_{\mathrm{evap}}-\tau)\frac{m_{H}r_{S}}{4\pi\hbar}\int^{\infty}_{l}\frac{d\sigma}{\sigma^{2}}\right)\\ &\Rightarrow\frac{-1}{4\pi\hbar}\frac{m_{H}\,r_{S}}{l}.\end{split}$ (27)
Here, the integral’s upper limit at $\sigma\rightarrow\infty$ naturally vanishes due to $1/\infty=0$, ensuring that only the boundary at $l$ contributes to the evaporation rate.
Since the evaporation rate is, mathematically, negative and non-zero, the transfer of states from the black hole to the radiation background is a dissipative, irreversible process under semi-classical assumptions. This follows from the fact that, in this case, a quantum state escaping the horizon cannot be reabsorbed by the black hole. Upon instantaneous measurement, the black hole’s instantaneous radius $l$ is equivalent to the Schwarzschild radius $r_{S}$. Substituting this into Eq. (27) and incorporating it into Eq. (18), we define the Lagrangian as a reduced relativistic term with the remaining torque contribution:
$E=\frac{1}{2}m_{H}-2\pi\frac{dL}{d\tau}.$ (28)
2.5 Hawking Mass $m_{H}$
Solving for the mass of the emitted Hawking particle $m_{H}$ will help determine the torque contribution. To determine the mass, recall Eq. (15): $m_{H}r^{2}\dot{\phi}=q\hbar$. We can use our conditions of azimuthal symmetry ($\phi=2\pi$) and proper time variation in $r$ to yield
@@ 2.5 Hawking Mass $m_{H}$
implying Eq. (13). Our obtained mass depends on the energy state $q$, whose evaporation rate relies strictly on the one-particle occupancy limit. This reinforces this mass relation as the mass of one Hawking particle, and we may gauge $q=1$ as a result. Conventionally, the Hawking particle mass is obtained by the reduced Compton mass formula, $m=\hbar/\lambda_{C}$. Since Hawking radiation originates at the horizon, the Compton wavelength corresponds to the black hole circumference, $\lambda_{C}=2\pi r_{S}$. This requires $q=1$ in Eq. (29); refer to footnote [^1] We eventually verify that $q=1$ is fundamental by evaluating Eq. (27) using the $q$-independent Compton mass definition of $m_{H}$.. From both viewpoints, we yield
@@
Furthermore, given a quantitatively fixed $q=1$ and provided $r_{S}m_{H}=\hbar/(2\pi)$ via Eq. (30), $L=q\hbar$ and $L=2\pi r_{S}m_{H}$ are both constants and equal. Therefore, either expression produces no torque when proper time-differentiated: $dL/d\tau=0$. This simplifies the energy of the Hawking particle as exactly the thermal energy of the black hole:
$E_{\mathrm{tot}}=\frac{\hbar}{8\pi GM}\equiv k_{B}T_{H}.$ (31)
@@ 2.6 Total Energy States $s$
${dq}=\frac{-1}{8\pi^{2}}\frac{d\tau}{l}.$ (32)
To relate this to the black hole’s evaporation, we impose the condition $d\tau=dl$, linking the change in proper time to the decrease in the black hole’s instantaneous radius. Since the Hawking particle follows a worldine with $ds=d\tau$, this establishes a direct correspondence between the emitted particle’s trajectory and the shrinking black hole radius.
For the instantaneous radius, the relevant range starts at the initial radius $r_{S,0}=2GM_{0}$ and ends at the smallest radius achievable: $r_{\text{QBH}}=2l_{P}$. Thus, integrating over the range $l\in[r_{S,0},2l_{P}]$, we obtain the following, with the negative sign flipping the integration limits:
@@
or equivantly $q=\ln\left({M_{0}}/{m_{P}}\right)/{8\pi^{2}}$ in terms of the black hole’s initial mass and the Planck mass. For black hole masses $0.000212\leq(M_{0}/M_{\odot})<4.13\times 10^{30}$ ($M_{\odot}=2\times 10^{30}$ kg is the solar mass), we find that $\lfloor q\rfloor=1$. Here, $\lfloor\bullet\rfloor$ represents the floor function, rounding down fractional values to the nearest whole integer; this is applied on the state number, as it ensures the physical logic of whole energy states in the context of quantum systems. For masses $M>4.13\times 10^{30}M_{\odot}\simeq 8.26\times 10^{60}\,\text{kg}$, we obtain $\lfloor q\rfloor>1$. To give a perspective, the mass of the observable universe is of order $10^{53}\sim 10^{54}$ kg (e.g. $\sim 6\times 10^{22}M_{\odot}=1.2\times 10^{53}$ kg in Ref. [19]).
For astrophysical black holes with a minimum mass of $3M_{\odot}$, a single quantum state is emitted each time a Hawking particle escapes the event horizon. The remaining quantum states on the black hole surface, therefore, correspond to the remaining number of surface particles along the event horizon. This defines the total number of quantum states as $\lfloor s\rfloor=N\lfloor q\rfloor$, where $N$ is given by the instantaneous mass $M$. Consequently, in terms of masses, the instantaneous total number of states is
$s=\frac{M^{2}}{8\pi^{2}m_{P}^{2}}\ln\left(\frac{M_{0}}{m_{P}}\right).$ (34)
Therefore, $\lfloor s_{0}\rfloor=N_{0}\lfloor q\rfloor$ is the initial total number of states.